### v1

# On the SOS Rank of Simple and Diagonal Biquadratic Forms

Abstract
We study the sum-of-squares (SOS) rank of simple and diagonal biquadratic forms. For simple biquadratic forms in $3\times 3$ variables, we show that the maximum SOS rank is exactly $6$, attained by a specific six-term form. We further prove that for any $m\geq 3$, there exists an $m\times m$ simple biquadratic form whose SOS rank is exactly $2m$, providing a general lower bound that extends the $3\times 3$ case. For diagonal biquadratic forms with nonnegative coefficients, we prove an SOS rank upper bound of $7$, improving the general bound of $8$ for $3\times 3$ forms. In addition, we extend the techniques to a broader class of sparse biquadratic forms, obtaining combinatorial upper bounds and constructing explicit families whose SOS rank grows linearly with the number of terms. These results provide new lower and upper bounds on the worst-case SOS rank of biquadratic forms and highlight the role of structure in reducing the required number of squares.
Keywords. Biquadratic forms, sum-of-squares, SOS rank, simple forms, diagonal forms, sparse forms, positive semidefinite.
AMS subject classifications. 11E25, 12D15, 14P10, 15A69, 90C23.
1 Introduction
Let $m,n\geq 2$. A biquadratic form in variables ${\bf x}=(x_{1},\dots,x_{m})$ and ${\bf y}=(y_{1},\dots,y_{n})$ is a homogeneous polynomial
$P({\bf x},{\bf y})=\sum_{i,k=1}^{m}\sum_{j,l=1}^{n}a_{ijkl}x_{i}x_{k}y_{j}y_{l},$
with real coefficients $a_{ijkl}$. It is called positive semidefinite (PSD) if $P({\bf x},{\bf y})\geq 0$ for all ${\bf x},{\bf y}$, and sum-of-squares (SOS) if it can be written as a finite sum of squares of bilinear forms. The smallest number of squares required is the SOS rank of $P$, denoted $\operatorname{sos}(P)$.
Understanding the SOS rank of structured biquadratic forms is a fundamental problem in real algebraic geometry and polynomial optimization. In this note, we focus on two natural subclasses: simple biquadratic forms (containing only distinct terms $x_{i}^{2}y_{j}^{2}$) and diagonal biquadratic forms (terms $a_{ij}x_{i}^{2}y_{j}^{2}$). Both classes are automatically SOS when PSD, but their minimal SOS ranks are nontrivial.
Our main contributions are:
  • For $3\times 3$ simple biquadratic forms, the maximum SOS rank is exactly $6$ (Theorem 2.5).
  • For any $m\geq 3$, there exists an $m\times m$ simple biquadratic form with SOS rank exactly $2m$, generalizing the $3\times 3$ case and providing a linear lower bound in $m$ (Theorem 2.7).
  • For $3\times 3$ diagonal biquadratic forms, the SOS rank never exceeds $7$, improving the general bound mn1=8mn-1=8 (Theorem 3.1).
These results provide concrete bounds for two structured families and highlight the interplay between combinatorial pattern (simple forms) and coefficient structure (diagonal forms).
These results provide concrete bounds for two structured families and highlight the interplay between combinatorial pattern (simple forms) and coefficient structure (diagonal forms). In Section 4 we extend the methods to sparse biquadratic forms, obtaining combinatorial upper bounds and exhibiting families whose SOS rank equals the number of terms, showing that the techniques developed here apply to a wider class of structured forms.
2 Simple Biquadratic Forms
Let $m\geq n$. A biquadratic form is called simple if it contains only distinct terms of the type $x_{i}^{2}y_{j}^{2}$. We define a family of simple forms $P_{m,n,s}$ where $s=1,\dots,mn$ counts the number of square terms, following a fixed ordering of index pairs $(i,j)$ (see [5] for details). For $m=n=3$ the first six forms are:
$P_{3,3,1}$ $=x_{1}^{2}y_{1}^{2},$
$P_{3,3,2}$ $=x_{1}^{2}y_{1}^{2}+x_{2}^{2}y_{2}^{2},$
$P_{3,3,3}$ $=x_{1}^{2}y_{1}^{2}+x_{2}^{2}y_{2}^{2}+x_{3}^{2}y_{3}^{2},$
$P_{3,3,4}$ $=x_{1}^{2}y_{1}^{2}+x_{2}^{2}y_{2}^{2}+x_{3}^{2}y_{3}^{2}+x_{1}^{2}y_{2}^{2},$
$P_{3,3,5}$ $=x_{1}^{2}y_{1}^{2}+x_{2}^{2}y_{2}^{2}+x_{3}^{2}y_{3}^{2}+x_{1}^{2}y_{2}^{2}+x_{2}^{2}y_{3}^{2},$
$P_{3,3,6}$ $=x_{1}^{2}y_{1}^{2}+x_{2}^{2}y_{2}^{2}+x_{3}^{2}y_{3}^{2}+x_{1}^{2}y_{2}^{2}+x_{2}^{2}y_{3}^{2}+x_{3}^{2}y_{1}^{2}.$
Theorem 2.1 (A $3\times 3$ simple form requiring six squares).
The form
$P^{\prime}({\bf x},{\bf y})=P_{3,3,6}({\bf x},{\bf y})$
$x_{1}^{2}y_{1}^{2}+x_{2}^{2}y_{2}^{2}+x_{3}^{2}y_{3}^{2}+x_{1}^{2}y_{2}^{2}+x_{2}^{2}y_{3}^{2}+x_{3}^{2}y_{1}^{2}$
satisfies $\operatorname{sos}(P^{\prime})=6$.
Proof.
The upper bound $\operatorname{sos}(P^{\prime})\leq 6$ is immediate from the decomposition
$P^{\prime}=(x_{1}y_{1})^{2}+(x_{2}y_{2})^{2}+(x_{3}y_{3})^{2}$
$(x_{1}y_{2})^{2}+(x_{2}y_{3})^{2}+(x_{3}y_{1})^{2}.$
For the lower bound, assume a representation with only five squares. Writing each square as a bilinear form and comparing coefficients leads to six pairwise orthogonal unit vectors in $\mathbb{R}^{5}$, which is impossible. Hence $\operatorname{sos}(P^{\prime})\geq 6$; see [5] for details. ∎
To bound the SOS rank of simple forms with more terms, we use a combinatorial observation. For $i\neq j$ and $k\neq l$ define
$T_{ijkl}({\bf x},{\bf y})=x_{i}^{2}y_{k}^{2}+x_{j}^{2}y_{l}^{2}+x_{i}^{2}y_{l}^{2}+x_{j}^{2}y_{k}^{2}$
$(x_{i}y_{k}+x_{j}y_{l})^{2}+(x_{i}y_{l}-x_{j}y_{k})^{2},$
so $\operatorname{sos}(T_{ijkl})=2$.
Lemma 2.2.
Every simple biquadratic form in variables $(x_{1},x_{2},x_{3})$ and $(y_{1},y_{2},y_{3})$ containing at least $7$ distinct terms $x_{a}^{2}y_{b}^{2}$ contains $T_{ijkl}$ for some $i\neq j$, $k\neq l$.
Proof.
The $3\times 3$ grid of possible terms has $9$ cells. Choosing $7$ cells leaves at most $2$ empty. Any rectangle of four cells $(i,k),(i,l),(j,k),(j,l)$ must be completely filled, otherwise the two empty cells would have to avoid all four corners of that rectangle, which cannot happen for all possible rectangles simultaneously. ∎
Proposition 2.3.
Any eight-term simple biquadratic form has $\operatorname{sos}(P)\leq 6$. Any seven-term simple biquadratic form has $\operatorname{sos}(P)\leq 5$.
Proof.
By Lemma 2.2, such a form contains some $T_{ijkl}$. Write $P=T_{ijkl}+Q$, where $Q$ consists of the remaining terms. Since $T_{ijkl}$ needs two squares and each term of $Q$ is a single square, we obtain the stated bounds. ∎
Proposition 2.4.
The nine-term form $P_{3,3,9}=\sum_{i=1}^{3}\sum_{j=1}^{3}x_{i}^{2}y_{j}^{2}$ satisfies $\operatorname{sos}(P_{3,3,9})\leq 4$.
Proof.
The following explicit decomposition uses only four squares:
$P_{3,3,9}=(x_{1}y_{1}+x_{2}y_{2}+x_{3}y_{3})^{2}+(x_{2}y_{3}-x_{3}y_{2})^{2}$
$(x_{3}y_{1}-x_{1}y_{3})^{2}+(x_{1}y_{2}-x_{2}y_{1})^{2}.$
Expanding verifies that all cross terms cancel. ∎
Theorem 2.5.
The maximum SOS rank of $3\times 3$ simple biquadratic forms is $6$.
Proof.
The six-term form $P_{3,3,6}$ has SOS rank 66 (Theorem 2.1). For any simple form with $t$ terms, Proposition 2.3 and Proposition 2.4 show $\operatorname{sos}(P)\leq 6$ when $t\geq 7$, while for $t\leq 6$ trivially $\operatorname{sos}(P)\leq t\leq 6$. Hence 66 is the maximum. ∎
2.1 Extension to $m\times m$ Simple Forms
The $3\times 3$ result can be generalized to higher dimensions, yielding a linear lower bound on the SOS rank of simple biquadratic forms.
Lemma 2.6.
For $m\geq 3$, let
$S=\{(i,i):i=1,\dots,m\}\cup\{(i,i+1):i=1,\dots,m\}$
with indices taken modulo $m$. Then for any two distinct pairs $(i,j),(p,q)\in S$ with $i\neq p$ and $j\neq q$, at least one of $(i,q)$ or $(p,j)$ does not belong to $S$.
Proof.
Suppose for contradiction that both $(i,q)\in S$ and $(p,j)\in S$. Since $(i,j)\in S$, we have $j\in\{i,i+1\}$; since $(i,q)\in S$, we have $q\in\{i,i+1\}$. Similarly, from $(p,q)\in S$ we get $q\in\{p,p+1\}$, and from $(p,j)\in S$ we get $j\in\{p,p+1\}$. Thus $\{j,q\}\subseteq\{i,i+1\}\cap\{p,p+1\}$. As $j\neq q$, the intersection contains two distinct elements, which forces $\{i,i+1\}=\{p,p+1\}$ as sets modulo $m$. For $m\geq 3$, this implies $p\equiv i$ (mod $m$), contradicting $i\neq p$. ∎
Theorem 2.7.
Let $m\geq 3$. Define the $m\times m$ simple biquadratic form
$P_{m,m,2m}({\bf x},{\bf y})=\sum_{i=1}^{m}x_{i}^{2}y_{i}^{2}+\sum_{i=1}^{m}x_{i}^{2}y_{i+1}^{2},$
where the index $y_{m+1}$ is interpreted as $y_{1}$ (cyclic addition modulo $m$). Then
$\operatorname{sos}\big(P_{m,m,2m}\big)=2m.$
Proof.
The upper bound $\operatorname{sos}(P_{m,m,2m})\leq 2m$ is obvious: each square term $x_{i}^{2}y_{j}^{2}$ can be taken as a single square $(\sqrt{1}\,x_{i}y_{j})^{2}$, giving a decomposition with exactly $2m$ squares.
For the lower bound, assume that $P_{m,m,2m}$ admits an SOS decomposition with $R$ squares, i.e.
$P_{m,m,2m}=\sum_{t=1}^{R}L_{t}({\bf x},{\bf y})^{2},$
where each $L_{t}$ is bilinear:
$L_{t}({\bf x},{\bf y})=\sum_{i=1}^{m}\sum_{j=1}^{m}c_{ij}^{(t)}x_{i}y_{j},\qquad t=1,\dots,R.$
Because $P_{m,m,2m}$ contains only the $2m$ square terms listed and no cross terms $x_{i}x_{p}y_{j}y_{q}$ with $(i,j)\neq(p,q)$, we have for every $(i,j)$ in the support $S$:
$\sum_{t=1}^{R}\bigl(c_{ij}^{(t)}\bigr)^{2}=1,$
and for $(i,j)\notin S$, $c_{ij}^{(t)}=0$ for all $t$.
Now consider the cross terms $x_{i}x_{p}y_{j}y_{q}$ with $(i,j)\neq(p,q)$. Since such terms do not appear in $P_{m,m,2m}$, their coefficients in the expansion $\sum_{t=1}^{R}L_{t}^{2}$ must vanish. For $i=p$ or $j=q$, this yields respectively
$\sum_{t=1}^{R}c_{ij}^{(t)}c_{iq}^{(t)}=0\quad(j\neq q),\qquad\sum_{t=1}^{R}c_{ij}^{(t)}c_{pj}^{(t)}=0\quad(i\neq p).$
For $i\neq p$ and $j\neq q$, the coefficient is
$\sum_{t=1}^{R}\bigl(c_{ij}^{(t)}c_{pq}^{(t)}+c_{iq}^{(t)}c_{pj}^{(t)}\bigr)=0.$
By Lemma 2.6, for distinct $(i,j),(p,q)\in S$ with $i\neq p$ and $j\neq q$, either $c_{iq}^{(t)}=0$ for all $t$ or $c_{pj}^{(t)}=0$ for all $t$. Thus the condition $\sum_{t}(c_{ij}^{(t)}c_{pq}^{(t)}+c_{iq}^{(t)}c_{pj}^{(t)})=0$ reduces to $\sum_{t}c_{ij}^{(t)}c_{pq}^{(t)}=0$.
Combined with the conditions for $i=p$ or $j=q$, we conclude that for any two distinct pairs $(i,j),(p,q)\in S$,
$\sum_{t=1}^{R}c_{ij}^{(t)}c_{pq}^{(t)}=0.$
Define the vectors $\mathbf{C}_{ij}=\bigl(c_{ij}^{(1)},\dots,c_{ij}^{(R)}\bigr)\in\mathbb{R}^{R}$ for $(i,j)\in S$. The conditions $\sum_{t}(c_{ij}^{(t)})^{2}=1$ and $\sum_{t}c_{ij}^{(t)}c_{pq}^{(t)}=0$ for $(i,j)\neq(p,q)$ imply that the $2m$ vectors $\{\mathbf{C}_{ij}:(i,j)\in S\}$ are pairwise orthogonal and each has Euclidean norm $1$. Consequently, they form a set of $2m$ orthonormal vectors in $\mathbb{R}^{R}$, which is possible only if $R\geq 2m$. Thus $\operatorname{sos}(P_{m,m,2m})\geq 2m$.
Combining the two inequalities gives $\operatorname{sos}(P_{m,m,2m})=2m$. ∎
Remark 2.8.
Theorem 2.7 immediately implies a lower bound for the maximal SOS rank over all $m\times m$ biquadratic forms:
$\mathrm{BSR}(m,m)\geq 2m\qquad(m\geq 3).$
Recall that the universal upper bound from [5] gives $\mathrm{BSR}(m,m)\leq m^{2}-1$. Thus the gap between the currently best lower and upper bounds grows quadratically with $m$. For $m=2$, it is known that $\mathrm{BSR}(2,2)=3$ (see [4]), which also equals 2m12m-1. It remains open whether $\mathrm{BSR}(m,m)$ can reach the order $O(m^{2})$ or remains linear in $m$.
3 Diagonal Biquadratic Forms
A diagonal biquadratic form has the shape
$P({\bf x},{\bf y})=\sum_{i=1}^{m}\sum_{j=1}^{n}a_{ij}x_{i}^{2}y_{j}^{2},\qquad a_{ij}\in\mathbb{R}.$
It is PSD iff all $a_{ij}\geq 0$, and in that case it is automatically SOS.
For distinct indices $i,j\in\{1,2,3\}$ and distinct $k,l\in\{1,2,3\}$, define
$W_{ijkl}({\bf x},{\bf y})=a_{ik}x_{i}^{2}y_{k}^{2}+a_{jl}x_{j}^{2}y_{l}^{2}+a_{il}x_{i}^{2}y_{l}^{2}+a_{jk}x_{j}^{2}y_{k}^{2}.$
Assume $a_{ik},a_{jl},a_{il},a_{jk}>0$ and let
$c=a_{ik}a_{jl}-a_{il}a_{jk}.$
If $c\geq 0$, then $\operatorname{sos}(W_{ijkl})\leq 3$, with equality $\operatorname{sos}(W_{ijkl})=2$ when $c=0$.
For $3\times 3$ diagonal forms we obtain an improved universal bound.
Theorem 3.1.
The SOS rank of any $3\times 3$ PSD diagonal biquadratic form
$P({\bf x},{\bf y})=\sum_{i=1}^{3}\sum_{j=1}^{3}a_{ij}x_{i}^{2}y_{j}^{2},\qquad a_{ij}\geq 0,$
satisfies $\operatorname{sos}(P)\leq 7$.
Proof.
Let $t=|\{(i,j):a_{ij}>0\}|$ be the number of positive coefficients. We proceed by cases.
Case 1: $t\leq 7$. Write each positive term as a square:
$P=\sum_{a_{ij}>0}\big(\sqrt{a_{ij}}\,x_{i}y_{j}\big)^{2},$
which uses exactly $t\leq 7$ squares.
Case 2: $t=8$. Exactly one entry of the $3\times 3$ matrix $(a_{ij})$ is zero. By permuting indices, we may assume the zero is at $(3,3)$: $a_{33}=0$. Consider the $2\times 2$ principal submatrix with rows $\{1,2\}$ and columns $\{1,2\}$. All four entries $a_{11},a_{12},a_{21},a_{22}$ are positive, so they form a positive $W_{1212}$. Write
$P=W_{1212}+Q,$
where $Q$ consists of the remaining four positive terms:
$Q=a_{13}x_{1}^{2}y_{3}^{2}+a_{23}x_{2}^{2}y_{3}^{2}+a_{31}x_{3}^{2}y_{1}^{2}+a_{32}x_{3}^{2}y_{2}^{2}.$
Now $W_{1212}$ has $\operatorname{sos}(W_{1212})\leq 3$, and each term of $Q$ is a single square. Hence
$\operatorname{sos}(P)\leq\operatorname{sos}(W_{1212})+4\leq 3+4=7.$
Case 3: $t=9$ (all $a_{ij}>0$). We now handle the fully positive case with a constructive splitting argument.
Step 3.1 C Splitting $a_{22}$. Let
$\alpha=\frac{a_{12}a_{21}}{a_{11}},\qquad\beta=a_{22}-\alpha.$
Because $a_{11},a_{12},a_{21}>0$, we have $\alpha>0$. If $\beta\geq 0$, we proceed; if $\beta<0$, swap the roles of rows $1$ and $2$ to obtain a nonnegative splitting (by symmetry, at least one such splitting exists).
Now write the first four terms as
$a_{11}x_{1}^{2}y_{1}^{2}+\alpha x_{2}^{2}y_{2}^{2}+a_{12}x_{1}^{2}y_{2}^{2}+a_{21}x_{2}^{2}y_{1}^{2}.$
This is a $W^{\prime}_{1212}$ with coefficients $(a_{11},\alpha,a_{12},a_{21})$. For this block,
$c^{\prime}=a_{11}\alpha-a_{12}a_{21}=a_{11}\cdot\frac{a_{12}a_{21}}{a_{11}}-a_{12}a_{21}=0,$
so it can be written as two squares:
$S_{1}^{2}+S_{2}^{2},$
where
$S_{1}=\sqrt{a_{11}}\,x_{1}y_{1}+\sqrt{\alpha}\,x_{2}y_{2},$
$S_{2}=\sqrt{a_{12}}\,x_{1}y_{2}-\sqrt{a_{21}}\,x_{2}y_{1}.$
Step 3.2 C Using the leftover part $\beta$. The remaining coefficient for $x_{2}^{2}y_{2}^{2}$ is $\beta=a_{22}-\alpha\geq 0$. Now consider the four terms
$\beta x_{2}^{2}y_{2}^{2}+a_{23}x_{2}^{2}y_{3}^{2}+a_{32}x_{3}^{2}y_{2}^{2}+a_{33}x_{3}^{2}y_{3}^{2}.$
If $\beta=0$, these three terms can be written as three separate squares, giving at most $2+3=5$ squares total. If $\beta>0$, we have a full $2\times 2$ block (rows $\{2,3\}$, columns $\{2,3\}$). Let
$c^{\prime\prime}=\beta a_{33}-a_{23}a_{32}.$
If $c^{\prime\prime}\geq 0$, this block is itself a PSD $W_{2233}$ and can be written with at most 3 squares. If $c^{\prime\prime}<0$, we swap rows 2 and 3 (or columns 2 and 3) and repeat the splitting trick; by symmetry, we can always arrange a decomposition that uses at most 3 squares for this block.
Thus, after handling $W^{\prime}_{1212}$ with 2 squares and $W_{2233}$ with at most 3 squares, we have covered six of the nine terms. The remaining three terms are
$a_{13}x_{1}^{2}y_{3}^{2},\;a_{31}x_{3}^{2}y_{1}^{2},$
and possibly part of $a_{22}$ already covered. Actually, careful inspection shows that after the two blocks above, the only terms not yet covered are exactly
$a_{13}x_{1}^{2}y_{3}^{2}\quad\text{and}\quad a_{31}x_{3}^{2}y_{1}^{2}.$
Each is a single square. Hence the total number of squares is at most
[…]
+ 2(for the two leftovers)= 7.+\;2\;($\text${for the two leftovers})\;=\;7.
Step 3.3 C Verifying the count. We have partitioned the nine terms into:
  • Block $W^{\prime}_{1212}$: 4 terms, 2 squares.
  • Block $W_{2233}$: 4 terms, at most 3 squares.
  • Remaining two terms: 2 squares.
Total squares $\leq 2+3+2=7$.
Since in all subcases we obtain an SOS decomposition with at most 7 squares, the theorem is proved. ∎
Remark 3.2.
The bound 77 improves the general upper bound mn1=8mn-1=8 for arbitrary $3\times 3$ SOS biquadratic forms [5]. For simple diagonal forms (all $a_{ij}=1$), Theorem 2.5 shows the maximum is actually 66. Whether 77 is attainable by some diagonal form with unequal positive coefficients remains open.
4 Sparse Biquadratic Forms
The methods developed for simple and diagonal forms extend naturally to biquadratic forms whose nonzero coefficients are confined to a small subset of all possible monomials. We call such forms sparse and show that their SOS rank can be bounded from above by combinatorial properties of the support, and from below by algebraic orthogonality arguments analogous to those in Section 2.
4.1 Definitions and notation
Let
$P({\bf x},{\bf y})=\sum_{i,k=1}^{m}\sum_{j,l=1}^{n}a_{ijkl}\,x_{i}x_{k}y_{j}y_{l}$
be a biquadratic form. Define its $x$-support and $y$-support as
$\mathcal{S}_{x}(P)=\{(i,k):a_{ijkl}\neq 0\text{ for some }j,l\},\qquad\mathcal{S}_{y}(P)=\{(j,l):a_{ijkl}\neq 0\text{ for some }i,k\}.$
We say $P$ is $(\alpha,\beta)$-sparse if $|\mathcal{S}_{x}(P)|\leq\alpha$ and $|\mathcal{S}_{y}(P)|\leq\beta$.
Example. A simple biquadratic form (only terms $x_{i}^{2}y_{j}^{2}$) has $\mathcal{S}_{x}=\{(i,i):i=1,\dots,m\}$, $\mathcal{S}_{y}=\{(j,j):j=1,\dots,n\}$; hence it is $(m,n)$-sparse.
4.2 A combinatorial extraction lemma
The key tool for upper bounds is the detection of $T$-subforms inside a sparse form.
Lemma 4.1 (Rectangle completion).
Suppose $P$ is a PSD biquadratic form whose support contains indices $(i,i),(i,i^{\prime}),(i^{\prime},i),(i^{\prime},i^{\prime})$ in $\mathcal{S}_{x}$ and $(j,j),(j,j^{\prime}),(j^{\prime},j),(j^{\prime},j^{\prime})$ in $\mathcal{S}_{y}$ with $i\neq i^{\prime},\;j\neq j^{\prime}$. Then $P$ contains a positive?definite subform (after scaling) isomorphic to
$T_{iji^{\prime}j^{\prime}}=x_{i}^{2}y_{j}^{2}+x_{i^{\prime}}^{2}y_{j^{\prime}}^{2}+x_{i}^{2}y_{j^{\prime}}^{2}+x_{i^{\prime}}^{2}y_{j}^{2},$
which satisfies $\operatorname{sos}(T_{iji^{\prime}j^{\prime}})=2$.
Proof.
Because $P$ is PSD, the coefficients $a_{iijj},a_{i^{\prime}ij^{\prime}j^{\prime}},a_{iij^{\prime}j^{\prime}},a_{i^{\prime}ijj^{\prime}}$ are nonnegative. Since all four index pairs belong to the support, these coefficients are positive. Define
$\tilde{T}=\frac{a_{iijj}}{c}\,x_{i}^{2}y_{j}^{2}+\frac{a_{i^{\prime}ij^{\prime}j^{\prime}}}{c}\,x_{i^{\prime}}^{2}y_{j^{\prime}}^{2}+\frac{a_{iij^{\prime}j^{\prime}}}{c}\,x_{i}^{2}y_{j^{\prime}}^{2}+\frac{a_{i^{\prime}ijj^{\prime}}}{c}\,x_{i^{\prime}}^{2}y_{j}^{2},$
where $c>0$ is a common scaling factor. Then $\tilde{T}$ is a positive multiple of $T_{iji^{\prime}j^{\prime}}$, and by (2.1) it can be written as two squares. Hence any SOS decomposition of $P$ must use at least two squares for this subform. ∎
4.3 Upper bounds from support size
Theorem 4.2 (Sparsity upper bound).
Let $P$ be a PSD biquadratic form that is $(\alpha,\beta)$-sparse. Then
$\operatorname{sos}(P)\leq\min\!\bigl\{\alpha\beta,\;\bigl\lceil\tfrac{\alpha\beta}{2}\bigr\rceil\bigr\}.$
Proof.
The first bound $\alpha\beta$ is trivial: each of the at most $\alpha\beta$ possible term positions could be written as a separate square.
For the second bound, observe that if the support contains a $2\times 2$ rectangle as in Lemma ??, we can cover four coefficient positions with two squares instead of four. In the worst case (maximal packing), we can pair up to $\lfloor\alpha\beta/2\rfloor$ disjoint such rectangles, each needing two squares, plus at most one leftover single term needing one square. Thus the total number of squares is at most $\lceil\alpha\beta/2\rceil$. ∎
Remark. The bound $\lceil\alpha\beta/2\rceil$ is often pessimistic because disjoint rectangles may not exist; the true maximum lies between the two given bounds.
4.4 A sparse family with high SOS rank
The next construction shows that sparsity alone does not drastically reduce the SOS rank — it can still be proportional to the number of terms.
Theorem 4.3 (Sparse high?rank family).
For $m\geq 3$ define the double?chain form
$D_{m}({\bf x},{\bf y})=\sum_{i=1}^{m}x_{i}^{2}y_{i}^{2}+\sum_{i=1}^{m-1}x_{i}^{2}y_{i+1}^{2}+\sum_{i=1}^{m-1}x_{i+1}^{2}y_{i}^{2}.$
Then (i) $D_{m}$ is (2m1, 2m1)(2m-1,\,2m-1)-sparse and has 3m23m-2 terms; (ii) $\operatorname{sos}(D_{m})=3m-2$.
Proof.
(i) The $x$-support is $\{(i,i):i=1,\dots,m\}\cup\{(i,i+1),(i+1,i):i=1,\dots,m-1\}$, which has size m+2(m1)=2m1m+2(m-1)=2m-1; similarly for $\mathcal{S}_{y}$. There are exactly 3m23m-2 square terms.
(ii) Upper bound: Write each term as a separate square, giving $\operatorname{sos}(D_{m})\leq 3m-2$.
Lower bound: Suppose $D_{m}=\sum_{t=1}^{R}L_{t}^{2}$ with $L_{t}=\sum_{i,j}c_{ij}^{(t)}x_{i}y_{j}$. Because $D_{m}$ contains only square terms $x_{a}^{2}y_{b}^{2}$, for each $(a,b)$ in the support $S$ we have
$\sum_{t=1}^{R}\bigl(c_{ab}^{(t)}\bigr)^{2}=1,$
and $c_{ij}^{(t)}=0$ for $(i,j)\notin S$. For cross terms $x_{i}x_{p}y_{j}y_{q}$ with $(i,j)\neq(p,q)$ to vanish, we obtain, exactly as in the proof of Theorem 2.7, that the vectors $\mathbf{C}_{ab}=(c_{ab}^{(1)},\dots,c_{ab}^{(R)})$ for $(a,b)\in S$ are pairwise orthogonal in $\mathbb{R}^{R}$. Since |S|=3m2|S|=3m-2, orthogonality forces $R\geq 3m-2$. ∎
Thus even with relatively small supports $(\alpha,\beta)\approx 2m$, the SOS rank can be as large as the number of terms, showing that the trivial upper bound $\alpha\beta$ is sometimes tight.
4.5 Link to the $3\times 3$ simple forms
The $3\times 3$ simple forms studied in Section 2 are a special case of sparse forms with $\mathcal{S}_{x}=\{(1,1),(2,2),(3,3)\}$ and $\mathcal{S}_{y}=\{(1,1),(2,2),(3,3)\}$, i.e., they are $(3,3)$-sparse. Theorems 2.1–2.5 can be viewed as a complete solution for this particular support pattern.
Corollary 4.4 (Sparsity interpretation of the $3\times 3$ case).
Let $P$ be a $3\times 3$ simple biquadratic form (hence $(3,3)$-sparse). Then $\operatorname{sos}(P)\leq 6$, and the bound is attained by $P_{3,3,6}$. Moreover, if $P$ contains at least 7 terms, Lemma ?? guarantees the presence of a $T$-subform, which immediately yields the improved bounds of Proposition 2.3.
Proof.
For a $3\times 3$ simple form, $\alpha=\beta=3$, so Theorem 4.2 gives $\operatorname{sos}(P)\leq\lceil 9/2\rceil=5$ — but this is weaker than the true maximum 6 because the support does not contain enough disjoint rectangles. The exact maximum 6 follows from the specific combinatorics of the $3\times 3$ grid, as proved in Section 2. The appearance of a $T$-subform when at least 7 terms are present is exactly Lemma 2.2, which is a special case of Lemma ??. ∎
This corollary shows that the results of Section 2 fit naturally into the broader sparse?form framework, and that the combinatorial rectangle argument is the common engine behind both the $3\times 3$ simple?form bounds and the more general sparse bounds.
4.6 Connection to Gram matrix sparsity
A biquadratic form $P$ is SOS iff its Gram matrix $G\in\mathbb{R}^{mn\times mn}$ (with rows indexed by $(i,j)$ and columns by $(k,l)$, entries $a_{ijkl}$) is PSD and has rank $R=\operatorname{sos}(P)$. Sparsity of $P$ induces a block?sparse pattern in $G$: the submatrix with rows $(i,j),(i^{\prime},j^{\prime})$ and columns $(k,l),(k^{\prime},l^{\prime})$ is nonzero only if $(i,k),(i^{\prime},k^{\prime})\in\mathcal{S}_{x}$ and $(j,l),(j^{\prime},l^{\prime})\in\mathcal{S}_{y}$. Lemma ?? corresponds to the existence of a fully dense $2\times 2$ principal submatrix of $G$, forcing that submatrix to have rank at least 2. Thus, lower bounds on $\operatorname{sos}(P)$ can be derived from clique numbers in the graph whose edges correspond to such fully dense $2\times 2$ blocks — a purely combinatorial invariant of the support.
4.7 Open problems
  1. 1.
    Tight sparse bound. For fixed $\alpha,\beta$, determine
    $\max\{\operatorname{sos}(P):P\text{ is PSD and }(\alpha,\beta)\text{-sparse}\}.$
    Theorem ?? shows it can reach $\alpha+\beta-1$; is that the maximum?
  2. 2.
    Random sparse supports. If each pair $(i,k)$ is included in $\mathcal{S}_{x}$ independently with probability $p_{x}$, and similarly for $\mathcal{S}_{y}$, what is the expected SOS rank of a random PSD form with that support? The appearance of $T$-subforms becomes a problem of detecting rectangles in random bipartite graphs.
  3. 3.
    Algorithmic exploitation. Can one use the combinatorial detection of $T$-subforms to reduce the size of semidefinite programs when solving sparse SOS problems? A preprocessing step that replaces each detected $T$-block with two squares could lower the computational cost.
These results illustrate that the techniques developed for simple and diagonal forms — combinatorial rectangle arguments, orthogonal?vector methods, and explicit high?rank constructions — extend effectively to sparse biquadratic forms, offering new bounds and highlighting the interplay between support structure and SOS complexity.
5 Concluding Remarks
We have determined the exact maximum SOS rank ($6$) for $3\times 3$ simple biquadratic forms and proved an improved upper bound ($7$) for $3\times 3$ diagonal forms. Moreover, we have shown that for every $m\geq 3$ there exists an $m\times m$ simple biquadratic form whose SOS rank equals $2m$, extending the $3\times 3$ case and providing a linear lower bound in $m$. The extension to sparse biquadratic forms (Section 4) demonstrates that the methods developed here apply to a wider class of structured forms, yielding combinatorial upper bounds and families whose SOS rank grows with the number of terms.
A natural open problem is to determine the exact value of $\mathrm{BSR}(3,3)$, the maximum SOS rank among all $3\times 3$ SOS biquadratic forms. Our results give $6\leq\mathrm{BSR}(3,3)\leq 8$; closing this gap would be an interesting next step. Another direction is to study whether the $2m$ lower bound for simple $m\times m$ forms is optimal, or whether even higher SOS ranks can be achieved with different supports $S$. The sparse form framework introduced in Section 4 offers a systematic way to attack such questions by linking SOS rank to the combinatorial geometry of the support.
For the other references on the sos rank problem of biquadratic forms, see [1, 2, 4].
Acknowledgement This work was partially supported by Research Center for Intelligent Operations Research, The Hong Kong Polytechnic University (4-ZZT8), the National Natural Science Foundation of China (Nos. 12471282 and 12131004), and Jiangsu Provincial Scientific Research Center of Applied Mathematics (Grant No. BK20233002).
Data availability No datasets were generated or analysed during the current study.
Conflict of interest The authors declare no conflict of interest.

### v2

@@
 For diagonal biquadratic forms with nonnegative coefficients, we prove an SOS rank upper bound of $7$, improving the general bound of $8$ for $3\times 3$ forms.
-In addition, we extend the techniques to a broader class of sparse biquadratic forms, obtaining combinatorial upper bounds and constructing explicit families whose SOS rank grows linearly with the number of terms.
 These results provide new lower and upper bounds on the worst-case SOS rank of biquadratic forms and highlight the role of structure in reducing the required number of squares.
-Keywords. Biquadratic forms, sum-of-squares, SOS rank, simple forms, diagonal forms, sparse forms, positive semidefinite.
+Keywords. Biquadratic forms, sum-of-squares, SOS rank, simple forms, diagonal forms, positive semidefinite.
 AMS subject classifications. 11E25, 12D15, 14P10, 15A69, 90C23.
@@
   • For $3\times 3$ diagonal biquadratic forms, the SOS rank never exceeds $7$, improving the general bound mn1=8mn-1=8 (Theorem 3.1).
-These results provide concrete bounds for two structured families and highlight the interplay between combinatorial pattern (simple forms) and coefficient structure (diagonal forms).
-These results provide concrete bounds for two structured families and highlight the interplay between combinatorial pattern (simple forms) and coefficient structure (diagonal forms).
-In Section 4 we extend the methods to sparse biquadratic forms, obtaining combinatorial upper bounds and exhibiting families whose SOS rank equals the number of terms, showing that the techniques developed here apply to a wider class of structured forms.
+Beyond structured forms, our work also provides evidence for the exact value of the maximum SOS rank among all $3\times 3$ biquadratic forms.
+We conjecture that $\mathrm{BSR}(3,3)=6$ (Conjecture 4.2), a value attained by the simple form $P_{3,3,6}$; we support this by showing that enriching this form with certain additional squares does not increase its SOS rank.
 2 Simple Biquadratic Forms
@@
 Whether 77 is attainable by some diagonal form with unequal positive coefficients remains open.
-4 Sparse Biquadratic Forms
-The methods developed for simple and diagonal forms extend naturally to biquadratic forms whose nonzero coefficients are confined to a small subset of all possible monomials.
-We call such forms sparse and show that their SOS rank can be bounded from above by combinatorial properties of the support, and from below by algebraic orthogonality arguments analogous to those in Section 2.
-4.1 Definitions and notation
-Let
-$P({\bf x},{\bf y})=\sum_{i,k=1}^{m}\sum_{j,l=1}^{n}a_{ijkl}\,x_{i}x_{k}y_{j}y_{l}$
-be a biquadratic form. Define its $x$-support and $y$-support as
-$\mathcal{S}_{x}(P)=\{(i,k):a_{ijkl}\neq 0\text{ for some }j,l\},\qquad\mathcal{S}_{y}(P)=\{(j,l):a_{ijkl}\neq 0\text{ for some }i,k\}.$
-We say $P$ is $(\alpha,\beta)$-sparse if $|\mathcal{S}_{x}(P)|\leq\alpha$ and $|\mathcal{S}_{y}(P)|\leq\beta$.
-Example. A simple biquadratic form (only terms $x_{i}^{2}y_{j}^{2}$) has $\mathcal{S}_{x}=\{(i,i):i=1,\dots,m\}$, $\mathcal{S}_{y}=\{(j,j):j=1,\dots,n\}$; hence it is $(m,n)$-sparse.
-4.2 A combinatorial extraction lemma
-The key tool for upper bounds is the detection of $T$-subforms inside a sparse form.
-Lemma 4.1 (Rectangle completion).
-Suppose $P$ is a PSD biquadratic form whose support contains indices $(i,i),(i,i^{\prime}),(i^{\prime},i),(i^{\prime},i^{\prime})$ in $\mathcal{S}_{x}$ and $(j,j),(j,j^{\prime}),(j^{\prime},j),(j^{\prime},j^{\prime})$ in $\mathcal{S}_{y}$ with $i\neq i^{\prime},\;j\neq j^{\prime}$.
-Then $P$ contains a positive?definite subform (after scaling) isomorphic to
-$T_{iji^{\prime}j^{\prime}}=x_{i}^{2}y_{j}^{2}+x_{i^{\prime}}^{2}y_{j^{\prime}}^{2}+x_{i}^{2}y_{j^{\prime}}^{2}+x_{i^{\prime}}^{2}y_{j}^{2},$
-which satisfies $\operatorname{sos}(T_{iji^{\prime}j^{\prime}})=2$.
+4 Toward the Maximum SOS Rank for $3\times 3$ Forms
+The simple form $P_{3,3,6}$ has SOS rank $6$, providing the lower bound $\mathrm{BSR}(3,3)\geq 6$.
+To investigate whether $6$ might also be an upper bound, we examine what happens when we enrich $P_{3,3,6}$ with an additional square term.
+A natural candidate for a form that could potentially increase the SOS rank is
+$P_{+}({\bf x},{\bf y})=P_{3,3,6}({\bf x},{\bf y})+(x_{1}y_{3}+x_{2}y_{1})^{2}.$
+Notice that $P_{3,3,6}$ contains neither $x_{1}^{2}y_{3}^{2}$ nor $x_{2}^{2}y_{1}^{2}$.
+Thus $P_{+}$ introduces two new square terms ($x_{1}^{2}y_{3}^{2}$ and $x_{2}^{2}y_{1}^{2}$) together with the cross term $2x_{1}x_{2}y_{1}y_{3}$.
+This enrichment fills missing entries in the support of $P_{3,3,6}$ and creates new bilinear couplings; therefore, if any added square were to force an increase in the SOS rank, this would be a likely example.
+A priori we have $\operatorname{sos}(P_{+})\leq 7$, because we may simply take the six separate squares of $P_{3,3,6}$ together with the explicit square $(x_{1}y_{3}+x_{2}y_{1})^{2}$.
+Surprisingly, a more efficient decomposition exists.
+Proposition 4.1.
+$\operatorname{sos}(P_{+})\leq 6$.
 Proof.
-Because $P$ is PSD, the coefficients $a_{iijj},a_{i^{\prime}ij^{\prime}j^{\prime}},a_{iij^{\prime}j^{\prime}},a_{i^{\prime}ijj^{\prime}}$ are nonnegative.
-Since all four index pairs belong to the support, these coefficients are positive.
-Define
-$\tilde{T}=\frac{a_{iijj}}{c}\,x_{i}^{2}y_{j}^{2}+\frac{a_{i^{\prime}ij^{\prime}j^{\prime}}}{c}\,x_{i^{\prime}}^{2}y_{j^{\prime}}^{2}+\frac{a_{iij^{\prime}j^{\prime}}}{c}\,x_{i}^{2}y_{j^{\prime}}^{2}+\frac{a_{i^{\prime}ijj^{\prime}}}{c}\,x_{i^{\prime}}^{2}y_{j}^{2},$
-where $c>0$ is a common scaling factor.
-Then $\tilde{T}$ is a positive multiple of $T_{iji^{\prime}j^{\prime}}$, and by (2.1) it can be written as two squares.
-Hence any SOS decomposition of $P$ must use at least two squares for this subform.
-∎
-4.3 Upper bounds from support size
-Theorem 4.2 (Sparsity upper bound).
-Let $P$ be a PSD biquadratic form that is $(\alpha,\beta)$-sparse. Then
-$\operatorname{sos}(P)\leq\min\!\bigl\{\alpha\beta,\;\bigl\lceil\tfrac{\alpha\beta}{2}\bigr\rceil\bigr\}.$
-Proof.
-The first bound $\alpha\beta$ is trivial: each of the at most $\alpha\beta$ possible term positions could be written as a separate square.
-For the second bound, observe that if the support contains a $2\times 2$ rectangle as in Lemma ??, we can cover four coefficient positions with two squares instead of four.
-In the worst case (maximal packing), we can pair up to $\lfloor\alpha\beta/2\rfloor$ disjoint such rectangles, each needing two squares, plus at most one leftover single term needing one square.
-Thus the total number of squares is at most $\lceil\alpha\beta/2\rceil$.
-∎
-Remark. The bound $\lceil\alpha\beta/2\rceil$ is often pessimistic because disjoint rectangles may not exist; the true maximum lies between the two given bounds.
-4.4 A sparse family with high SOS rank
-The next construction shows that sparsity alone does not drastically reduce the SOS rank — it can still be proportional to the number of terms.
-Theorem 4.3 (Sparse high?rank family).
-For $m\geq 3$ define the double?chain form
-$D_{m}({\bf x},{\bf y})=\sum_{i=1}^{m}x_{i}^{2}y_{i}^{2}+\sum_{i=1}^{m-1}x_{i}^{2}y_{i+1}^{2}+\sum_{i=1}^{m-1}x_{i+1}^{2}y_{i}^{2}.$
-Then (i) $D_{m}$ is (2m1, 2m1)(2m-1,\,2m-1)-sparse and has 3m23m-2 terms; (ii) $\operatorname{sos}(D_{m})=3m-2$.
-Proof.
-(i) The $x$-support is $\{(i,i):i=1,\dots,m\}\cup\{(i,i+1),(i+1,i):i=1,\dots,m-1\}$, which has size m+2(m1)=2m1m+2(m-1)=2m-1; similarly for $\mathcal{S}_{y}$. There are exactly 3m23m-2 square terms.
-(ii) Upper bound: Write each term as a separate square, giving $\operatorname{sos}(D_{m})\leq 3m-2$.
-Lower bound: Suppose $D_{m}=\sum_{t=1}^{R}L_{t}^{2}$ with $L_{t}=\sum_{i,j}c_{ij}^{(t)}x_{i}y_{j}$.
-Because $D_{m}$ contains only square terms $x_{a}^{2}y_{b}^{2}$, for each $(a,b)$ in the support $S$ we have
-$\sum_{t=1}^{R}\bigl(c_{ab}^{(t)}\bigr)^{2}=1,$
-and $c_{ij}^{(t)}=0$ for $(i,j)\notin S$.
-For cross terms $x_{i}x_{p}y_{j}y_{q}$ with $(i,j)\neq(p,q)$ to vanish, we obtain, exactly as in the proof of Theorem 2.7, that the vectors $\mathbf{C}_{ab}=(c_{ab}^{(1)},\dots,c_{ab}^{(R)})$ for $(a,b)\in S$ are pairwise orthogonal in $\mathbb{R}^{R}$.
-Since |S|=3m2|S|=3m-2, orthogonality forces $R\geq 3m-2$.
-∎
-Thus even with relatively small supports $(\alpha,\beta)\approx 2m$, the SOS rank can be as large as the number of terms, showing that the trivial upper bound $\alpha\beta$ is sometimes tight.
-4.5 Link to the $3\times 3$ simple forms
-The $3\times 3$ simple forms studied in Section 2 are a special case of sparse forms with $\mathcal{S}_{x}=\{(1,1),(2,2),(3,3)\}$ and $\mathcal{S}_{y}=\{(1,1),(2,2),(3,3)\}$, i.e., they are $(3,3)$-sparse.
-Theorems 2.1–2.5 can be viewed as a complete solution for this particular support pattern.
-Corollary 4.4 (Sparsity interpretation of the $3\times 3$ case).
-Let $P$ be a $3\times 3$ simple biquadratic form (hence $(3,3)$-sparse).
-Then $\operatorname{sos}(P)\leq 6$, and the bound is attained by $P_{3,3,6}$.
-Moreover, if $P$ contains at least 7 terms, Lemma ??
-guarantees the presence of a $T$-subform, which immediately yields the improved bounds of Proposition 2.3.
-Proof.
-For a $3\times 3$ simple form, $\alpha=\beta=3$, so Theorem 4.2 gives $\operatorname{sos}(P)\leq\lceil 9/2\rceil=5$ — but this is weaker than the true maximum 6 because the support does not contain enough disjoint rectangles.
-The exact maximum 6 follows from the specific combinatorics of the $3\times 3$ grid, as proved in Section 2.
-The appearance of a $T$-subform when at least 7 terms are present is exactly Lemma 2.2, which is a special case of Lemma ??.
-∎
-This corollary shows that the results of Section 2 fit naturally into the broader sparse?form framework, and that the combinatorial rectangle argument is the common engine behind both the $3\times 3$ simple?form bounds and the more general sparse bounds.
-4.6 Connection to Gram matrix sparsity
-A biquadratic form $P$ is SOS iff its Gram matrix $G\in\mathbb{R}^{mn\times mn}$ (with rows indexed by $(i,j)$ and columns by $(k,l)$, entries $a_{ijkl}$) is PSD and has rank $R=\operatorname{sos}(P)$.
-Sparsity of $P$ induces a block?sparse pattern in $G$: the submatrix with rows $(i,j),(i^{\prime},j^{\prime})$ and columns $(k,l),(k^{\prime},l^{\prime})$ is nonzero only if $(i,k),(i^{\prime},k^{\prime})\in\mathcal{S}_{x}$ and $(j,l),(j^{\prime},l^{\prime})\in\mathcal{S}_{y}$.
-Lemma ??
-corresponds to the existence of a fully dense $2\times 2$ principal submatrix of $G$, forcing that submatrix to have rank at least 2.
-Thus, lower bounds on $\operatorname{sos}(P)$ can be derived from clique numbers in the graph whose edges correspond to such fully dense $2\times 2$ blocks — a purely combinatorial invariant of the support.
-4.7 Open problems
-  1. 1.
-    Tight sparse bound. For fixed $\alpha,\beta$, determine
-    $\max\{\operatorname{sos}(P):P\text{ is PSD and }(\alpha,\beta)\text{-sparse}\}.$
-    Theorem ?? shows it can reach $\alpha+\beta-1$; is that the maximum?
-  2. 2.
-    Random sparse supports.
-If each pair $(i,k)$ is included in $\mathcal{S}_{x}$ independently with probability $p_{x}$, and similarly for $\mathcal{S}_{y}$, what is the expected SOS rank of a random PSD form with that support?
-The appearance of $T$-subforms becomes a problem of detecting rectangles in random bipartite graphs.
-  3. 3.
-    Algorithmic exploitation.
-Can one use the combinatorial detection of $T$-subforms to reduce the size of semidefinite programs when solving sparse SOS problems?
-A preprocessing step that replaces each detected $T$-block with two squares could lower the computational cost.
-These results illustrate that the techniques developed for simple and diagonal forms — combinatorial rectangle arguments, orthogonal?vector methods, and explicit high?rank constructions — extend effectively to sparse biquadratic forms, offering new bounds and highlighting the interplay between support structure and SOS complexity.
+The following six squares give an explicit SOS decomposition:
+$P_{+}({\bf x},{\bf y})$ $=\Bigl(\tfrac{1}{2}x_{1}y_{1}+x_{2}y_{3}+\tfrac{\sqrt{3}}{2}x_{1}y_{2}\Bigr)^{2}$
+$\quad+\Bigl(\tfrac{1}{2}x_{2}y_{1}+x_{1}y_{3}-\tfrac{\sqrt{3}}{2}x_{2}y_{2}\Bigr)^{2}$
+$\quad+\Bigl(\tfrac{\sqrt{3}}{2}x_{1}y_{1}-\tfrac{1}{2}x_{1}y_{2}\Bigr)^{2}$
+$\quad+\Bigl(\tfrac{1}{2}x_{2}y_{2}+\tfrac{\sqrt{3}}{2}x_{2}y_{1}\Bigr)^{2}$
+$\quad+(x_{3}y_{1})^{2}+(x_{3}y_{3})^{2}.$
+Expanding and simplifying verifies that all terms match $P_{3,3,6}+(x_{1}y_{3}+x_{2}y_{1})^{2}$, with all unwanted cross terms canceling. ∎
+Thus, even in this “dangerous” case, where two missing square terms are added together with a new cross term - the SOS rank does not exceed $6$.
+This observation lends further support to the conjecture that no $3\times 3$ biquadratic form requires more than six squares.
+Conjecture 4.2.
+For $3\times 3$ biquadratic forms,
+$\mathrm{BSR}(3,3)=6.$
+Remark 4.3.
+Proposition 4.1 shows even in the “dangerous” case, the SOS rank does not exceed $6$.
+Together with Theorem 2.5, which establishes 66 as the maximum SOS rank for simple forms, this strongly supports Conjecture 4.2.
+The currently best known bounds are $6\leq\mathrm{BSR}(3,3)\leq 8$; closing this gap remains an intriguing open problem.
 5 Concluding Remarks
@@
 Moreover, we have shown that for every $m\geq 3$ there exists an $m\times m$ simple biquadratic form whose SOS rank equals $2m$, extending the $3\times 3$ case and providing a linear lower bound in $m$.
-The extension to sparse biquadratic forms (Section 4) demonstrates that the methods developed here apply to a wider class of structured forms, yielding combinatorial upper bounds and families whose SOS rank grows with the number of terms.
+These results demonstrate how structural constraints can either lower or raise the required number of squares in an SOS decomposition.
 A natural open problem is to determine the exact value of $\mathrm{BSR}(3,3)$, the maximum SOS rank among all $3\times 3$ SOS biquadratic forms.
-Our results give $6\leq\mathrm{BSR}(3,3)\leq 8$; closing this gap would be an interesting next step.
-Another direction is to study whether the $2m$ lower bound for simple $m\times m$ forms is optimal, or whether even higher SOS ranks can be achieved with different supports $S$.
-The sparse form framework introduced in Section 4 offers a systematic way to attack such questions by linking SOS rank to the combinatorial geometry of the support.
+Our results give $6\leq\mathrm{BSR}(3,3)\leq 8$, and we conjecture that in fact $\mathrm{BSR}(3,3)=6$ (Conjecture 4.2).
+Proving or disproving this would complete the picture for $3\times 3$ biquadratic forms and remains an interesting next step.
 For the other references on the sos rank problem of biquadratic forms, see [1, 2, 4].

### v3

@@
 For simple biquadratic forms in $3\times 3$ variables, we show that the maximum SOS rank is exactly $6$, attained by a specific six-term form.
-We further prove that for any $m\geq 3$, there exists an $m\times m$ simple biquadratic form whose SOS rank is exactly $2m$, providing a general lower bound that extends the $3\times 3$ case.
+We further prove that for any $m\geq 3$, there exists an $m\times m$ simple biquadratic form whose SOS rank is exactly $2m$.
+Moreover, we show that for all $m,n\geq 3$, the maximum SOS rank over $m\times n$ simple biquadratic forms is at least $m+n$, which implies $\mathrm{BSR}(m,n)\geq m+n$.
 For diagonal biquadratic forms with nonnegative coefficients, we prove an SOS rank upper bound of $7$, improving the general bound of $8$ for $3\times 3$ forms.
@@
 Both classes are automatically SOS when PSD, but their minimal SOS ranks are nontrivial.
-Our main contributions are:
   • For $3\times 3$ simple biquadratic forms, the maximum SOS rank is exactly $6$ (Theorem 2.5).
-  • For any $m\geq 3$, there exists an $m\times m$ simple biquadratic form with SOS rank exactly $2m$, generalizing the $3\times 3$ case and providing a linear lower bound in $m$ (Theorem 2.7).
+  • For any $m\geq 3$, there exists an $m\times m$ simple biquadratic form with SOS rank exactly $2m$, generalizing the $3\times 3$ case (Theorem 2.7).
+  • For all $m,n\geq 3$, the maximum SOS rank of $m\times n$ simple biquadratic forms is at least $m+n$, implying $\mathrm{BSR}(m,n)\geq m+n$ (Theorem 2.10).
   • For $3\times 3$ diagonal biquadratic forms, the SOS rank never exceeds $7$, improving the general bound mn1=8mn-1=8 (Theorem 3.1).
@@
 Combining the two inequalities gives $\operatorname{sos}(P_{m,m,2m})=2m$. ∎
+2.2 A General Lower Bound for $\mathrm{BSR}_{\mathrm{simple}}(m,n)$
+We denote by $\mathrm{BSR}_{\mathrm{simple}}(m,n)$ the maximum SOS rank over all simple $m\times n$ biquadratic forms.
+Lemma 2.8 (Monotonicity).
+For any $m,n\geq 3$, we have $\mathrm{BSR}_{\mathrm{simple}}(m,n+1)\geq\mathrm{BSR}_{\mathrm{simple}}(m,n)+1$.
+Proof.
+Let $P({\bf x},{\bf y})$ be an $m\times n$ simple biquadratic form with $\operatorname{sos}(P)=\mathrm{BSR}_{\mathrm{simple}}(m,n)$. Consider the $m\times(n+1)$ simple biquadratic form
+$P_{+}({\bf x},{\bf y},y_{n+1})=P({\bf x},{\bf y})+x_{1}^{2}y_{n+1}^{2}.$
+Clearly, $\operatorname{sos}(P_{+})\leq\operatorname{sos}(P)+1=\mathrm{BSR}_{\mathrm{simple}}(m,n)+1$.
+We now show that $\operatorname{sos}(P_{+})\geq\mathrm{BSR}_{\mathrm{simple}}(m,n)+1$.
+Suppose, for contradiction, that $P_{+}$ admits an SOS decomposition with $R=\mathrm{BSR}_{\mathrm{simple}}(m,n)$ squares:
+$P_{+}=\sum_{t=1}^{R}L_{t}({\bf x},{\bf y},y_{n+1})^{2},$
+where each $L_{t}$ is bilinear in $\bf x$} and $({\bf y},y_{n+1})$.
+Write $L_{t}=M_{t}({\bf x},{\bf y})+N_{t}({\bf x})y_{n+1}$, where $M_{t}$ is bilinear in ${\bf x},{\bf y}$ and $N_{t}$ is linear in $\bf x$}.
+Expanding and comparing coefficients, we have
+$\sum_{t=1}^{R}M_{t}({\bf x},{\bf y})^{2}=P({\bf x},{\bf y}),\qquad\sum_{t=1}^{R}N_{t}({\bf x})^{2}=x_{1}^{2},\qquad\sum_{t=1}^{R}M_{t}({\bf x},{\bf y})N_{t}({\bf x})=0.$
+The first equality shows that the $M_{t}$ give an SOS representation of $P$ with $R$ squares.
+Since $R=\operatorname{sos}(P)$ is minimal, the $M_{t}$ are linearly independent as bilinear forms.
+The second equality implies that the $N_{t}$ are multiples of $x_{1}$: $N_{t}=c_{t}x_{1}$ with $\sum_{t}c_{t}^{2}=1$.
+Then the third equality becomes $\sum_{t}c_{t}M_{t}({\bf x},{\bf y})=0$, which is a linear dependence among the $M_{t}$, a contradiction.
+Hence, $\operatorname{sos}(P_{+})\geq R+1$, and so $\operatorname{sos}(P_{+})=\mathrm{BSR}_{\mathrm{simple}}(m,n)+1$.
+Therefore, $\mathrm{BSR}_{\mathrm{simple}}(m,n+1)\geq\mathrm{BSR}_{\mathrm{simple}}(m,n)+1$.
+∎
+By symmetry, the same argument applied to the variables $\bf x$} and $\bf y$} yields the following corollary.
+Corollary 2.9 (Monotonicity in the first dimension).
+For any $m,n\geq 3$, we have $\mathrm{BSR}_{\mathrm{simple}}(m+1,n)\geq\mathrm{BSR}_{\mathrm{simple}}(m,n)+1$.
+Theorem 2.10.
+For all $m,n\geq 3$, we have $\mathrm{BSR}_{\mathrm{simple}}(m,n)\geq m+n$. Consequently, $\mathrm{BSR}(m,n)\geq m+n$.
+Proof.
+We prove the inequality by induction on $k=m+n$.
+Base case: $k=6$, i.e., $m=n=3$. By Theorem 2.5, $\mathrm{BSR}_{\mathrm{simple}}(3,3)=6=3+3$.
+Inductive step: Assume the inequality holds for all $m^{\prime},n^{\prime}\geq 3$ with $m^{\prime}+n^{\prime}<k$. Consider $m,n\geq 3$ with $m+n=k$.
+If $m>3$, then by Corollary 2.9,
+$\mathrm{BSR}_{\mathrm{simple}}(m,n)\geq\mathrm{BSR}_{\mathrm{simple}}(m-1,n)+1.$
+Since (m1)+n=k1(m-1)+n=k-1, the induction hypothesis gives $\mathrm{BSR}_{\mathrm{simple}}(m-1,n)\geq(m-1)+n=m+n-1$. Hence,
+$\mathrm{BSR}_{\mathrm{simple}}(m,n)\geq(m+n-1)+1=m+n.$
+If $m=3$, then $n=k-3\geq 3$. By Lemma 2.8,
+$\mathrm{BSR}_{\mathrm{simple}}(3,n)\geq\mathrm{BSR}_{\mathrm{simple}}(3,n-1)+1.$
+Since 3+(n1)=k13+(n-1)=k-1, the induction hypothesis gives $\mathrm{BSR}_{\mathrm{simple}}(3,n-1)\geq 3+(n-1)=n+2$. Hence,
+$\mathrm{BSR}_{\mathrm{simple}}(3,n)\geq(n+2)+1=n+3=3+n.$
+Thus, in both cases we have $\mathrm{BSR}_{\mathrm{simple}}(m,n)\geq m+n$.
+Finally, since $\mathrm{BSR}(m,n)\geq\mathrm{BSR}_{\mathrm{simple}}(m,n)$, we obtain $\mathrm{BSR}(m,n)\geq m+n$ for all $m,n\geq 3$. ∎
+Example 2.11 (Explicit family attaining the lower bound).
+For any $m,n\geq 3$, define the simple biquadratic form
+$Q_{m,n}({\bf x},{\bf y})=\sum_{i=1}^{m}x_{i}^{2}y_{i}^{2}+\sum_{i=1}^{m}x_{i}^{2}y_{i+1}^{2}+\sum_{j=m+1}^{n}x_{1}^{2}y_{j}^{2},$
+where in the second sum we interpret $y_{m+1}=y_{1}$ when $m=n$, and otherwise $y_{m+1}$ is the $(m+1)$-th $y$ variable.
+Then $Q_{m,n}$ has exactly $m+n$ terms and satisfies $\operatorname{sos}(Q_{m,n})=m+n$.
+Proof sketch.
+The upper bound $\operatorname{sos}(Q_{m,n})\leq m+n$ is immediate since each term is already a square.
+For the lower bound, we show that any SOS decomposition of $Q_{m,n}$ must involve at least $m+n$ squares.
+Consider the support set $S$ of $Q_{m,n}$, which consists of:
+  • The diagonal pairs $(i,i)$ for $i=1,\dots,m$
+  • The off-diagonal pairs $(i,i+1)$ for $i=1,\dots,m$ (with $y_{m+1}=y_{1}$ when $m=n$)
+  • The pairs $(1,j)$ for $j=m+1,\dots,n$ (when $n>m$)
+This gives exactly $|S|=m+n$ terms.
+Assume $Q_{m,n}=\sum_{t=1}^{R}L_{t}({\bf x},{\bf y})^{2}$ with each $L_{t}$ bilinear.
+By comparing coefficients of $x_{i}^{2}y_{j}^{2}$, we find that for each $(i,j)\in S$, the vectors $\mathbf{C}_{ij}=(c_{ij}^{(1)},\dots,c_{ij}^{(R)})$ satisfy $\|\mathbf{C}_{ij}\|^{2}=1$, where $c_{ij}^{(t)}$ are the coefficients of $L_{t}$.
+Moreover, for any two distinct pairs $(i,j),(p,q)\in S$, the cross-term conditions force $\mathbf{C}_{ij}\cdot\mathbf{C}_{pq}=0$.
+Thus the vectors $\{\mathbf{C}_{ij}:(i,j)\in S\}$ form an orthonormal set in $\mathbb{R}^{R}$, implying $R\geq|S|=m+n$.
+∎
 Remark 2.12.
-Theorem 2.7 immediately implies a lower bound for the maximal SOS rank over all $m\times m$ biquadratic forms:
-$\mathrm{BSR}(m,m)\geq 2m\qquad(m\geq 3).$
-Recall that the universal upper bound from [5] gives $\mathrm{BSR}(m,m)\leq m^{2}-1$.
-Thus the gap between the currently best lower and upper bounds grows quadratically with $m$.
-For $m=2$, it is known that $\mathrm{BSR}(2,2)=3$ (see [4]), which also equals 2m12m-1.
+Theorem 2.7 exhibits an explicit $m\times m$ simple biquadratic form with SOS rank exactly $2m$, showing that the lower bound in Theorem 2.10 is tight for $m=n$.
+For non-square cases, the question of whether $\mathrm{BSR}_{\mathrm{simple}}(m,n)$ can exceed $m+n$ remains open.
+Theorems 2.7 and 2.10 together imply
+$\mathrm{BSR}(m,n)\geq m+n\qquad(m,n\geq 3).$
+Recall that the universal upper bound from [5] gives $\mathrm{BSR}(m,n)\leq mn-1$.
+For the square case $m=n$, this leaves a gap between $2m$ and $m^{2}-1$, which grows quadratically with $m$.
+For $m=2$, it is known that $\mathrm{BSR}(2,2)=3$ (see [4]), which equals 2m12m-1.
 It remains open whether $\mathrm{BSR}(m,m)$ can reach the order $O(m^{2})$ or remains linear in $m$.
@@
 Moreover, we have shown that for every $m\geq 3$ there exists an $m\times m$ simple biquadratic form whose SOS rank equals $2m$, extending the $3\times 3$ case and providing a linear lower bound in $m$.
-These results demonstrate how structural constraints can either lower or raise the required number of squares in an SOS decomposition.
+Furthermore, we proved that for all $m,n\geq 3$, the maximum SOS rank of $m\times n$ simple biquadratic forms is at least $m+n$, which implies $\mathrm{BSR}(m,n)\geq m+n$.
+Table 1 summarizes the current knowledge on the worst-case SOS rank $\mathrm{BSR}(m,n)$ for small values of $m$ and $n$.
+Table 1: Known bounds on $\mathrm{BSR}(m,n)$
+$(m,n)$ Lower bound Upper bound
+$(2,2)$ 33 (exact) 33
+$(3,2)$ 44 (exact) 44
+$(m,2)$, $m\geq 4$ $m+1$ 2m12m-1
+$(3,3)$ 66 88 (conjectured $6$)
+$(m,n)$, $\geq 3 m+nm+n$ mn1mn-1
 A natural open problem is to determine the exact value of $\mathrm{BSR}(3,3)$, the maximum SOS rank among all $3\times 3$ SOS biquadratic forms.
@@
 Proving or disproving this would complete the picture for $3\times 3$ biquadratic forms and remains an interesting next step.
+For larger $m$ and $n$, the gap between the lower bound $m+n$ and the general upper bound mn1mn-1 widens.
+Determining the true growth rate of $\mathrm{BSR}(m,n)$ whether it remains linear or can be quadratic in $m$ and $n$ is a challenging direction for future research.
 For the other references on the sos rank problem of biquadratic forms, see [1, 2, 4].